Notation

Let U denote a representative for a conjugacy class of subgroups of the group G with multiplicity-free permutation character π. The information for each conjugacy class consists of four parts:

1. The isomorphism type of U: the notation U ≤ H means that U is contained in the maximal subgroup H of G, and if no ≤ sign occurs then U is itself maximal in G,

2. The degree π(1) = |G|/|U|,

3. The decomposition of π into irreducible characters of G, and

4. The rank of π.